find all finite abelian group of order less than or equal to 40 up to isomorphism.
1
Expert's answer
2020-05-24T19:01:49-0400

Trivial group "G_1 = \{e\}" is abelian group of order 1.

Groups of prime order less than 40 are cyclic and hence abelian, there is one and only one group upto isomorphism per prime "p, Z_p" where "p" equal to "2,3,5,7,11,13,17,19,23,29,31,37".

Divide the composite number into its prime factor, such as factors gcd not equal to one. For example divide "Z_{20}" as "Z_2\times Z_{10}" both are abelian group of same order.

However if gcd is 1, then they are isomorphic. For example: "Z_4\times Z_5 \cong Z_{20}"


List of all abelian group of order less than or equal to 40 utpo isomorphism are:

"Z_1 \\
Z_2 \\
Z_3 \\
Z_4 \\
Z_2 \times Z_2 \\ 
Z_5 \\
Z_6 \\
Z_7 \\
Z_8 \\
Z_4 \times Z_2 \\
Z_2 \times Z_2 \times Z_2 \\
Z_9 \\
Z_3 \times Z_3 \\
Z_{10} \\
Z_{11} \\
Z_{12} \\
Z_6 \times Z_2 \\
Z_{13} \\
Z_{14} \\
Z_{15} \\
Z_{16} \\
Z_4 \times Z_4 \\
Z_8 \times Z_2 \\
Z_4 \times Z_2 \times Z_2 \\
Z_2 \times Z_2 \times Z_2 \times Z_2 \\
Z_{17} \\
Z_{18} \\
Z_6 \times Z_3 \\
Z_{19} \\
Z_{20} \\
Z_{10} \times Z_2 \\
Z_{21} \\
Z_{22} \\
Z_{23} \\
Z_{24} \\ 
Z_{12} \times Z_2 \\
Z_6 \times Z_2 \times Z_2 \\
Z_{25} \\
Z_5 \times Z_5 \\
Z_{26} \\
Z_{27} \\
Z_9 \times Z_3 \\
Z_3 \times Z_3 \times Z_3 \\
Z_{28} \\
Z_{14} \times Z_2 \\
Z_{29} \\
Z_{30} \\
Z_{31} \\
Z_{32} \\
Z_8 \times Z_4 \\
Z_{16} \times Z_2 \\
Z_4 \times Z_4 \times Z_2 \\
Z_8 \times Z_2 \times Z_2 \\
Z_4 \times Z_2 \times Z_2 \times Z_2 \\
Z_2 \times Z_2 \times Z_2 \times Z_2 \times Z_2 \\
Z_{33} \\
Z_{34} \\
Z_{35} \\
Z_{36} \\
Z_{18} \times Z_2 \\
Z_{12} \times Z_3 \\
Z_6 \times Z_6 \\
Z_{37} \\
Z_{38} \\
Z_{39}\\
Z_{40} \\
Z_{20} \times Z_2 \\
Z_{10} \times Z_2 \times Z_2."